Kendra has $4.20 in dimes and nickels. If she has 3 times as many dimes as nickels, how many of each coin does she have?
Answer by richwmiller(17219)

5n+10d=420
n+2d=84
3n=d
3n-d=0
n+2d=84
3,-1,0
1.2,0
(12n,36d)